OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

–x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (anonymous):

@heyitsmelalala You can use the point-slope form to know the equation. \[(y-y _{1})=m(x-x _{1})\] \[(y-4)=(-1)(x+1)\] Distribute -1. \[y-4=-x-1\] Transpose -4: \[y=-x+3\] You may say, "Those aren't on the choices" but in fact, they are. You just need to twist those choices and you'll see that the answer is: \[x+y=3\]

OpenStudy (anonymous):

Ohh okay thank you! I get it now

OpenStudy (anonymous):

@heyitsmelalala You're welcome. :)